OpenStudy (anonymous):

convert r=2a(1+sin(2t)) into cartesian equation

OpenStudy (anonymous):

nbd, sin(2t)=2sin(t)cos(t)=2(y/r)(x/r)=2yx/r^2

OpenStudy (anonymous):

whats nbd and how do u 2sin(t)cos(t)=sin(2t) is that an identity?

hartnn (hartnn):

The formulae i am gonna use : sin(2t)=2sint*cost 1=sint^2+cost^2 then sint=y/r.cost=x/r,and r=root (x^2+y^2)  i skipped few steps,because of space....but tell me if u did not get it...